Assume that the company has the following capital structure:

Debt

$15,000,000

Preferred stock

$7,500,000

Common stock

$27,500,000

What will be the cost of capital if the company decide to raise the needed capital proportionally and with following costs? Please use the following information to calculate the weighted cost of capital:

  1. Bond:

A 30-year bond with a face value of $1000 and coupon interest rate of 13% and floatation cost of $20 (Tax is 35%)

  1. Preferred stock:

Face value of $35 that pays dividend $5 and floatation cost of $2

  1. Common stock:

Market value of $54 with floatation cost of $3.5. Last dividend was $6. The dividend will expect to grow at 7%.
Answer #1

Answer:

Capital structure (including weights of individual components) is:

Cost of equity = (Dividend next year / (Share price - Flotation cost)) + dividend growth rate

= (6 * (1 + 7%) / (54 - 3.50)) + 7%

=19.71%

Cost of Preferred stock = Dividend / (Face value - Flotation cost))

= 5 / (35 - 2)

= 15.15%

Cost of debt:

Annual coupon = 1000 * 13% = $130

Time to maturity = 30 years

Issue proceeds = 1000 - 20 = $980

Hence:

Cost of debt = RATE (nper, pmt, pv, fv, type)

= RATE (30, 130, -980, 1000, 0)

= 13.27%

WACC:

WACC = Cost of equity * weight of equity + Cost of preferred stock * weight of preferred stock + Cost of debt * (1 - Tax rate) * weight of debt

= 19.71% * 55% + 15.15% * 15% + 13.27% * (1 - 35%) * 30%

= 15.70%

Cost of capital = 15.70%